OpenStudy (pphalke):

Find the total monthly payment for a $275,000 mortgage loan at 5.5% for 30 years. The assessed value of the home is $300,000. The annual taxes on the home are 1.2% of the assessed value, and the insurance on the home costs $750 per year

OpenStudy (wolf1728):

Here is the monthly payment loan formula http://www.1728.org/loanform.htm Months = 30*12 = 360 Annual Taxes = .012 * 300,000 = 3,600 times 30 = TOTAL TAXES = 108,000 TOTAL INSURANCE = 750/year * 30 years = 22,500 So, before we figure out the mortgage, do we agree that the total mortgage cost would be 275,000 plus 108,000 plus 22,500 405500 TOTAL